KTeV DAY Shift Summary - 09/05/99 (Sun.)

Earl C. Swallow (earls@elmhurst.edu)
Sun, 5 Sep 1999 15:23:47 -0500 (CDT)

Shift Summary for DAY shift, 09/05/99 (Sun.)

On Shift: Elizabeth Turner & Earl Swallow + EC and other visitors

Hours of beam available at all intensities: 7.3 (out of 8.5)
Hours of beam available at nominal intensity: 7.3
Hours of E832 data (special runs not included): 7.1
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time & intensity:

1 tape change & several KFRENDS & some too-high intensity
+ 1 big quench & accelerator energy conservation & aborts.

========================================================================
Runs ended this shift:
14430 - BEAM_832

Tapes brought to FCC this shift:
QK(C,S,K)387-389, QKB136 (Run 14428)
QK(C,S,K)390-392, QKB137 (Runs 14429[muon], 14430)

========================================================================
Accesses during shift:

None, of course.

========================================================================
Overview of shift:

Quiet shift with a big quench and some too-high intensity episodes,
but generally steady data taking.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14430 07:00 Shift Begins.
Tapes are almost full. All else is well.
The ignorable intermittant VolMon2 error remains.

14430 07:15 END Run 14430 - BEAM_832.
Remove tapes: QK(C,S,K)390-392, QKB137.

14431 1 07:20 BEGIN Run 14431 - BEAM_832.
Mounted Tapes: QK(C,S,K)393-395, QKB138. NOTE:
Again, both members of Shift Crew checked tape
numbers before eating lunch.

08:49 TeVatron quenches at LOTS of locations. It could
take some time to recover - est. 45 minutes.

87 09:31 Beam Returns -- with 2 calls by 2 different
operators from MCR -- well under estimate!

10:00 Beam intensity is a bit high with some intermittant
time structure in the 0.2-1 Hz creating chamber
current excursions up to ~150 muA, but it does appear
to be slowly improving with time.

11:20 Intensity lower and more stable (consistently below
6.95E12) with most chamber currents in the
85-125 muA region and few DAQ complaints.

186 11:45 Intensity up as high as 7.3E12 -- call MCR.

12:15 Filled out Gas Check List, and just after the
Ar-Ethane had switched from Center to West, DN
appears to change bottles and check on system. The
gas system for this run has worked like a Swiss
watch -- BRAVO!
However, we are out of blank check list forms --
email sent requesting more and two auxillary rows
added at bottom of completed form.

14:40 P. McBride comes through with more Gas Check List
forms. They are now in the front of Gas Check List
notebook.

1431 15:30 Shift ends with continued data taking.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

EB says to do a Laser Scan if there is appropriate down time.

--
===============================================================================
Earl C. Swallow                           email:  earls@elmhurst.edu
Theophil W. Mueller Professor              (or)   earls@fnal.gov
   and Chairperson
Department of Physics                     Phone:  (630)-617-3577
Elmhurst College                          FAX:    (630)-617-3735
190 Prospect Avenue
Elmhurst, IL   60126-3296                 http://www.elmhurst.edu/~earls
===============================================================================
"Books are lighthouses erected in the dark sea of time."
                                --  from Gargoyles  (The Scrolls of Merlin)